Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis 

Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 74,000 units of product were as follows: 


Standard CostsActual Costs
Direct materials251,600 lbs. at $5.40249,100 lbs. at $5.20
Direct labor18,500 hrs. at $17.3018,930 hrs. at $17.70
Factory overheadRates per direct labor hr.,

based on 100% of normal

capacity of 19,310 direct

labor hrs.:

Variable cost, $4.80$87,910 variable cost

Fixed cost, $7.60$146,756 fixed cost


Each unit requires 0.25 hour of direct labor. 


Required: 

a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using minus sign and an unfavorable variance as a positive number.

Direct Materials Price Variance

Direct Materials Quantity Variance

Total Direct Materials Cost Variance


b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sig and an unfavorable variance as a positive number.

Direct Labor Rate Variance

Direct Labor Time Variance

Total Direct Labor Cost Variance


c. Determine the variable factory overhead controllable variance, fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

 Variable factory overhead controllable variance 

 Fixed factory overhead volume variance 

 Total factory overhead cost variance
Answer #1
Variance analysis is a technical jargon to explain a situation where actual result differs from planned
or expected results. It is an act of comparing standards with actual. The study of variances help in
decision making and finding out the major areas which needs immediate attention.
MAckinaw Inc.
Variances
MPV = ( SP - AP ) * AQ purchased
MPV = ( 5.40 - 5.2 ) * 249100 = $ 49820 ( favourable )
MQV = ( SQ - AQ ) SP
MQV = ( 251600 - 249100 ) * 5.40 = $ 13500 ( favourable )
Total DM cost variance = MPV + MQV = 49820 + 13500 = 63320 ( favourable )
LRV = ( SR - AR ) AH
LRV = ( 17.3 - 17.7 ) 18930 = $ 7572 (UN- Favourable)
LEV = ( SH - AH ) SR
LEV = ( 18500 - 18930 ) 17.3 = $ 7439 ( unfavourable)
Total DL cost variance = LRV + LEV = 7572 + 7439 = 15011 ( unfavourable )
VOH controllable varaince = Standard cost allowed - Actual variable OH cost
74000 * 0.25 * 4.80 - 87910 = $ 890 ( favourable )
FOH Volume Variance = ( SH for Actual output - Budgeted hours ) * SR
( 74000 * 0.25 - 19310 ) * 7.60 = $ 6156 ( favourable )
Total factory OH cost variance = VOH controllable variance + FOH volume variance
890 + 6156 = 7046 ( favourable )